Need help solving for x: rx + h = 5x - k

Answer by stanbon(75887)

rx + h = 5x - k
-------------------
rx - 5x = -(k+h)
----
x(r-5) = (k+h)
-----
x = (k+h)/(r-5)
-------------------
Cheers,
Stan H.
----------------